OpenStudy (anonymous):

Find the first partial derivative of the function. f(r,s)=r*ln(r^2+s^2)

OpenStudy (valpey):

\[f(r,s) = r*\ln(g(r,s))\]\[\frac{\partial{f(r,s)}}{\partial{r}} = \ln(g(r,s))+r*\frac{\frac{\partial{g(r,s)}}{\partial{r}}}{g(r,s)}\]

OpenStudy (lgbasallote):

why couldnt it have been for s lol

OpenStudy (anonymous):

I need r and s!

OpenStudy (lgbasallote):

first find the derivative of r (treat s as a constant) what do you get?